OpenStudy (anonymous):

what is the center of the following ellipse.. x^2 (y+2)^2 ____ + _______ =1 1 9

OpenStudy (anonymous):

\((0,-2)\) since it looks like \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] with \(h=0,k=-2\)

OpenStudy (anonymous):

thats what i thought also but then they trough a (-2,0) in there and i got confused on what came first thank you :)

OpenStudy (anonymous):

yw